Operations Report – January 14th

Operations Report

Subject Line: Crew 187 Operations Report 14JAN2018

Crew 187 Operations Report 14012018

SOL: 2

Name of person filing report: L. Diaz

Non-nominal systems: None

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15h 50min

Generator turned off, charging battery at 9:45 am

Generator turned on at 18:36 pm

Solar— SOC

@ 11:04: 100%

@ 18:36: 53%

Diesel: 50%

Propane: 27%

Ethanol Free Gasoline (5 Gallon containers for ATV): XX Gallons

Water (trailer): – Gallons

Water (static): – Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: – Gallons

Toilet tank emptied: No

ATVs Used: (Honda, 350.1, 350.2, 350.3)

Oil Added? No

ATV Fuel Used: – Gallons

# Hours the ATVs were used today: 0.5 hours

Notes on ATVs: ATVs were nominal.

Deimos rover used: No

Hours: –

Beginning charge: 100%

Ending charge:

Currently charging: No

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge:

Currently charging: Maybe

Spirit rover used: Si

Hours: –

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours: –

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: –

Beginning charge: 100%

Ending charge:

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Today was just a training day. I didn’t have the opportunity to check the hours from the rover’s, fuel and the level of water.

Summary of internet: All nominal

Summary of suits and radios: All nominal but I found that the ventilation system of the regular suit #5 is not working.

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: As I said in the Summary of suits, the ventilation system of the regular suit #5 is not working, can I try to fix it tomorrow?

Journalist Report – January 14th

[Sol 02] [Interstellar]

The day began with the farewell and boarding of the Crew 186 in their spaceship to our planet and old sweet home, Earth. From this dawn we became Martians coming from distant lands between each other: Peru, Colombia and Mexico, but in this desert planet we only represent the Earth and some of its forms of life.

We grew up thinking that astronauts only eat tubes of dehydrated food but on this trip to Mars we discovered that macaroni with cheese, tuna and chocolate cookies will also be part of our diet. Important: We receive food from the Russian Space Food Laboratory, thank you.

In our arrival to the habitat we have settled in our small rooms that barely reach a few square meters that daily will see us sleeping while thousands of stars shine in the sky and a body similar to the satellite of the Earth rises between the Martian mountains illuminating together with the Sun our mornings.

Today the habitat director showed us every corner of the station and the operation. Also we use of Spirit, Curiosity, Deimos, the Red ATV and the three Blue 350 vehicles to explore some roads and nearby lands. We also were showed the Science Lab and the solar observatory for in the next few days be able to use them.

When we returned to the habitat we made our first pressurization test where for five minutes we locked ourselves in the airlock to discuss which songs will be the ones that relax us during the minutes that the process will take. Apparently telling jokes will also be allowed.

Finally the night fell and after writing the reports and planning the activities of the next day we will sit down to eat pancakes and food experiments as well as continue knowing the pleasures and expectations of each crew member.

We hope to start our first EVA tomorrow morning and use our suits along with the helmets to train but not without first performing a morning exercise routine and taking a nutritious breakfast.

LATAM II will continue to inform.

Tania Robles, MDRS Crew 187 Journalist

GreenHab Report – January 14th

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:

Ambient with door opening


Shade cloth on

Working Hour: 6:30
Inside temp at working hour: 17° C
Outside temp during working hours: 0° C
Inside temperature high: 30° C
Inside temperature low: 15 C
Inside humidity: 75 %RH

Inside humidity high: 75 %RH
Inside humidity low: 21 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: No changes

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 15 min

Research observations: The research has not started

Changes to research plants: The research has not started

Aquaponics: Not Functional

Narrative: Today I have started as the new Greenhab Officer in the station, I hope to continue maintaining it as beautiful as now. Thank you Mark.

During my first day, I have divided the crops in 8 regions in order to compare each day the changes using picture

Tomorrow I will start to assemble the lisimiter in order to start to measure evapotranspiration in the crop of quinoa that Atila plant at the Saturday. Also, I will check the inventory of the crops that Mark shared with me during the training.

Today during the irrigation, I realized that one of the sprinklers is broken.

Sol Summary – January 14th

Crew 187 Sol 2 Summary Report 14JAN2018

Sol 2

Summary Title: Across the Universe

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Mission Support,

Greetings from Mars!!! After a very long journey we finally arrived to Mars yesterday, where we spent some quality time with the previous crew. Today is our first day managing everything by ourselves, we awoke early at 06:30 for the farewell of crew 186. We are so excited to be in Mars, we started our day talking about the rules and the schedules that we are going to follow during this awesome experience. Then we settle down in our home. One of the surprise of the day was the moment when we opened the food suppliers and found food from Russian Space Food Laboratory… Thank you so much for it!

For lunch Oscar cooked us tuna salad and potato… our first lunch in Mars!!! By the afternoon, Dr. Ruppert explain us in detail the protocols that we have to follow and we had our EVA’s training with her. Later, we started to plan the following EVAs.

Look Ahead Plan:

Dr. Ruppert will tell us the final part of the protocols of the simulation in the morning. Later we are going to do 2 EVAs in order to assets the performance of the crew with in suits operations. It’s almost some hours more and we are going to know the feeling of walking and be a part of expedition in Mars.

Anomalies in work:

We suspect that there is a leak in the toilet. The water level’s start to increasing little by little but then it works normal again. We hope it continuing working in a good way.

Weather: Temperature: 4°C, Fair weather, Winds calm, Humidity 60%, Barometer 30.35 inHg

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, Operations Report, Journalist Report, Greenhab Report, EVA Request

Support Requested: Mission Support, please add my email to the e-list of MDRS Mission Support.

Copyright © The Mars Society. All rights reserved. | Main Site